Consider 5 independent Bernoulli’s trials each with probability of success p. If the probability of at least
one failure is greater than or equal to
31 /32
, then p lies in the interv

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
n=5
success=p
failue=q
P (at least one failure)\geq(31/32)
1 – P (no failure)\geq(31/32)
1- P(x=5)\geq(31/32)
1 5C5p5\geq(31/32)
on simplifying we get
p\leq(1/2)
so p\epsilon(0,1/2)
Thanks and Regards
Shaik Aasif
askIITians faculty